An applied force of 109 N is used to accelerate an object to the right across a frictional surface. The object encounters 19.4 N of friction. The object weighs 88.5 N.

a) Calculate the normal force,

(b) calculate the net force,

(,c) calculate the coeficcient of friction (mu, µ),

(d) calculate the mass

(e) calculate the acceleration.

please write neatly when explaining thank you
Answer #1

Solution)

Part a) Normal force N= 88.5 N

======

Part b) Net force=109 N- 19.4 N = 89.6 N

======

Part c) coefficient of friction, mu= f/N = 19.4/88.5 = 0.219

=======

Part d) here, 88.5= mg

So, m = 88.5/9.8 =9.02 kg

==========

Part e) acceleration a= net force/m = 89.6/ 9.02 = 9.83 m/s^2

========

Good luck!:)